# Multi-View Projection for Unsupervised Domain Adaptation in 3D Semantic Segmentation

### v2

@@ Abstract
3D semantic segmentation plays a pivotal role in autonomous driving and road infrastructure analysis, yet state-of-the-art 3D models are prone to severe domain shift when deployed across different datasets. In this paper, we propose a novel multi-view projection framework that excels in unsupervised domain adaptation (UDA). Our approach first aligns Lidar scans into coherent 3D scenes and renders them from multiple virtual camera poses to create large-scale synthetic 2D datasets (PC2D) in various modalities. We then use them to train an ensemble of 2D segmentation models in pointcloud view domain on each modality. During inference, the models process hundreds of views per scene; the resulting logits are back-projected to 3D with an occlusion-aware voting scheme to generate final point-wise labels. These labels can be used directly or to fine-tune a 3D segmentation model in the target domain. We evaluate our approach in both Real-to-Real and Simulation-to-Real UDA settings. We achieve state-of-the-art results in UDA on Real-to-Real. We also demonstrate how our method can be used to segment ‘rare classes’, for which target 3D annotations are not available, by only using 2D annotations for those classes and leveraging 3D annotations for other classes in a source domain.
I Introduction
@@
3D semantic segmentation of Lidar scenes is crucial for applications such as autonomous driving (AD) and road infrastructure analysis. In particular, many applications require models to be able to generalize to unseen data, which may not be from the same domain as the training data, e.g. if the data was acquired in a different location or with a different sensor. The main attempt to mitigate this problem is Unsupervised Domain Adaptation (UDA).
@@
Most existing methods are tailored for AD, thus making concessions on accuracy to guarantee real-time performance on single or few Lidar scans.
We propose a method that leverages aligned scenes and larger computation times to perform 3D semantic segmentation. In particular, it reaches state-of-the-art-level performance in UDA.
@@
Our method belongs to the multi-view projection family, where 3D scenes are analyzed via multiple 2D representations. We position virtual cameras around aligned Lidar scenes to generate rendered views from various perspectives. While similar methods use camera images for training and inference [1], we create large-scale synthetic datasets by rendering both the 3D scenes and their corresponding ground truth annotations. This enables a 2D segmentation model to be trained on data closely matching the inference domain, allowing to freely choose the optimal parameters for selecting and rendering the views. The extensive use of 2D models leads to interesting domain adaptation performance. We demonstrate the effectiveness of our method for unsupervised domain adaptation in Real-to-Real tasks on the nuScenes [2] and SemanticKITTI [3] datasets, and in a Simulation-to-Real task on the SynLidar [4] dataset. We also demonstrate a potential application of our method to segment ‘rare classes’, where 3D annotations are not available for the desired target classes, but 2D annotations are available along with 3D annotations for other classes. Finally, we conduct a full ablation study to verify the importance of each component of our method. Our main contributions are the following:
A novel modular multi-view projection framework with dataset generation capabilities that can be used to explore many approaches and parameters for 3D semantic segmentation
@@
A state-of-the-art method for unsupervised domain adaptation in 3D semantic segmentation
@@
We demonstrate an application of our method to segment ‘rare classes’ (defined above)
II Related works
II-A 3D semantic segmentation
@@ II Related works
3D semantic segmentation is the task of assigning a semantic label to each point in a 3D pointcloud. It is generally done either on Lidar pointclouds [2, 3], or RGB-D images [5]. Methods are usually divided in 3D-based and 2D-based. 3D models directly operate on the pointcloud using 3D sparse convolutions [6], while 2D models re-use 2D image segmentation methods by projecting the pointcloud to a 2D plane [7]. Importantly, 3D models have been shown to suffer from domain shift
@@
to a higher degree than 2D models [8]. 2D approaches generally project the pointcloud to a single 2D image with channels representing various features, e.g. height (in Bird’s Eye View), intensity, density, etc. Multi-view projection methods try to leverage 2D models even further by repeatedly using the same model on multiple views of the same 3D scene.
II-B Multi-view projection for segmentation
@@
1.
Which views to use for inference and with which model ?
2.
@@
Our proposed method uses a simple voting scheme on accumulated logits from a single 2D model.
In this paper, we focus on the first problem. Initially, [16] proposed multi-view projection for indoor RGB-D scenes, where the views are simply the RGB-D images themselves. For outdoor Lidar scenes, the same reasoning led [1] to use aligned RGB camera images as views, for which 2D annotated data is available [17]. However, this approach is limited as it requires using additional sensors, and the performance is tied to the quality of the sensor alignment. Mitigating this, [12] proposed to infer on rendered views of 3D scenes with an out-of-domain 2D model trained on camera images. This has the advantage of not requiring 3D labels or additional sensors, but comes with several limitations:
The domain gap between camera images and views of Lidar pointclouds is substantial, which negatively impacts 2D model performance.
The multi-view aspect of the method is limited because the view generation method must align with the camera images, e.g., be from a road user’s perspective.
@@
We propose to solve all of these limitations by leveraging 3D annotations. This makes our method more dependent on 3D annotated data, but allows to reach significantly better performance and to tackle dynamic classes. Closest to our approach, [10] proposes to use 3D annotations to generate 2D synthetic datasets to train a 2D model for downstream multi-view projection for indoor RGB-D scenes. Our method distinguishes itself in its application to outdoor Lidar scenes, where much additional processing is required to obtain satisfying views, e.g. scan aligning and meshing, as well as in the choices for the feature rendering and view selection. We also propose an ensembling mechanism for handling the modality choice, and demonstrate the potential of such methods for unsupervised domain adaptation.
II-C Unsupervised Domain Adaptation
Unsupervised Domain Adaptation (UDA) aims to adapt a model trained on a labeled source domain to perform well on an unlabeled target domain , where no ground-truth labels are available. Unlike supervised adaptation, only the source labels and unlabeled target samples can be exploited, while the evaluation is performed on a disjoint set . This setting is particularly relevant for autonomous driving, where annotated data is costly to acquire, but raw scans from the deployment domain can be collected at scale.
@@ II-C Unsupervised Domain Adaptation
Our method differs by relying entirely on a multi-view 2D projection pipeline to generate pseudo-labels for the target domain. These pseudo-labels are then used to fine-tune a 3D segmentation model. This fundamentally different approach allows to reach state-of-the-art performance in UDA, especially on large, structured classes.
III Method
III-A Base method
An overview of our method, called Seg_3D_by_PC2D , is shown in Figure 1. We choose to use [12] as a base 3D multi-view projection method, which we enhance and extend.
We provide a brief description of that method below. Here, a scene refers to the concatenation of multiple Lidar scans aligned in a common reference frame. Note that this concatenation yields denser scenes, but may introduce so called ‘shadow effects’ in the scene, due to dynamic classes moving during capture.
1.
Train a 2D model for 2D semantic segmentation on natural camera images.
2.
Render views of a 3D scenes, then process them with the model, which is out of domain.
3.
Project the 2D segmentations on the 3D pointcloud and accumulate the logits, as votes.
4.
For each point, assign the most voted label.
This method yields a 3D segmentation mask while using no 3D annotation at all. Our key innovation is to bridge the domain gap between camera images and rendered views of Lidar pointclouds by re-training the 2D model in pointcloud view domain, using 3D annotations. This allows to drastically improve the performance of the base method, and can still be applied without 3D annotations for the target classes, see the rare classes experiment IV-E2.
First, we use 3D data along with 3D annotations to generate large-scale synthetic datasets of views of 3D scenes along with 2D segmentation masks, as shown in Figure 2, which we use to train the 2D model. We call these datasets ‘PointCloud2D’ datasets (PC2D). The parameters of the view generation are chosen carefully, as they are the same that will be used for inference downstream. We optimize the number and type of virtual camera poses (VGO, III-C), as well as the visualization modalities used. We use multiple models trained on each modality, and use them together in as an ensemble (MMods, III-D). Additionally, we improve the base method’s projection by adding occlusion using depth maps (OCL, III-E). In the experiments, we validate all of these components in the ablation study IV-F.
III-B Training in pointcloud view domain using 3D annotations in 2D (PC2D)
In [12], there is a substantial gap between the domain of the generated views and the domain of the camera training images. The main extension we propose to the base method is to train the 2D model in a much closer domain using 3D annotations. To this end, we generate a large-scale synthetic dataset of 2D views of 3D scenes along with 2D segmentation masks, as shown in Figure 2.
@@ III-B Training in pointcloud view domain using 3D annotations in 2D (PC2D)
Virtual camera poses are sampled around the 3D scene, which are then used to render 2D views of the scene. This way, we can generate an arbitrary number of (image, label) pairs, yielding a large-scale synthetic dataset. Training on such a synthetic dataset can be seen as a powerful data augmentation. The only limitation is that a limited number of annotated scenes will lead to samples with less diversity as the total number of training samples increases, which could lead to overfitting.
@@
The use of PC2D datasets allows to choose the domain of the training images, which allows full freedom over the view generation steps during both training and inference e.g. generating bird’s eye view depth maps is possible, while views in [12], were limited to ‘car’ view rendering of Lidar intensity. The two following subsections describe how we optimize the camera pose generation and the modality choice to obtain the best performance.
III-C View Generation Optimization (VGO)
@@
We propose to optimize the view generation by choosing relevant camera poses. In [12], the camera poses are all placed at a road user’s perspective to match training data. We keep this category and add 3 new categories of camera poses. All 4 types of camera poses are obtained based on the original Lidar sensor poses, which are already describing the trajectory of the ego vehicle along the scene. An illustration of the 4 types of camera pose is shown in Figure 2, in the VGO module, and we describe them below:
@@
car_view: The camera is placed at a road user’s perspective, i.e. at the approximate height of a car (random yaw can be applied).
@@
conic_view: The camera is placed on the basis of a large imaginary cone pointing downwards, with the apex at the Lidar sensor position. The camera points to the apex of the cone. This corresponds to a ‘drone’ or ‘lamp post’ view of the scene, and is also the closest to the views that human annotators would typically use to annotate a scene.
@@
top_view: The camera is placed directly above the Lidar position, looking downwards, with variable height and roll. This corresponds to the usual ‘bird’s eye view’ of the scene.
@@
bottom_view: The camera is placed directly below the Lidar position, looking upwards. This is the opposite of the top view. Rendering Lidar pointclouds from below allows to clearly see the colors of the flat surfaces of the ground without occlusions from the objects above.
We use all 4 types of camera poses, with 250 camera poses per type, per scene, thus 1,000 total views per scene. Using more views seems not to improve performance, and using fewer views leads to a drop.
III-D Multi-Modalities (MMods)
@@
Similar to camera pose generation, PC2D datasets allow us to choose the visualization modality for the training and inference images. While RGB cameras are limited to RGB colors, we can choose to render modalities such as Lidar intensity, depth maps, and normals. The final projection and voting step III-E does not limit the number of votes, we therefore propose to train multiple models on multiple modalities, and use them together in an ensemble. We use the following modalities:
Intensities: The Lidar sensor intensity, normalized by scan.
RGB: AD Lidar datasets often provide RGB images, which can be used to colorize the pointclouds through projection. We render views of the obtained colored pointclouds.
Depths: The pointcloud depth maps with respect to the camera pose.
Normals: The components of the pointcloud surface normals, computed using the normal estimation method in [23].
For each modality, we generate a PC2D dataset, and train a separate model. We then use these 4 models to process each view of the scene at inference time, accumulating their segmentation outputs as votes. If a given dataset does not provide a modality (e.g. RGB), the remaining models can be used individually. The ablation study IV-F shows the performance is only slightly impacted when not using the RGB modality (-0.1 mIoU).
III-E Projection and Voting
@@
We use a similar projection step as in [12], where 2D logits are back-projected to 3D and accumulated as votes. We propose adding occlusion to the projection step. Naive back-projection from 2D to 3D suffers from points being included in the projection even if they are occluded by other points. We use depth maps generated from meshes, themselves generated using [23] for occlusion. Using meshes has the advantage of allowing less sparse depth maps which is more appropriate for dense segmentation masks. Finally, we set a margin parameter to include points behind the depth map to account for the size of the occluding object.
IV Experiments
| NSSK10 (% IoU) | Car | Bicycle | Motorcycle | Truck | Other vehicle | Pedestrian | Driveable surf. | Sidewalk | Terrain | Vegetation | % mIoU | |
| Source-only | 74.09 | 0.01 | 10.18 | 1.07 | 1.0 | 4.43 | 63.44 | 31.81 | 36 | 25.6 | 24.76 | |
| T-UDA | [18] | 93.0 | 0.0 | 11.4 | 3.4 | 47.0 | 15.7 | 83.3 | 54.4 | 67.9 | 83.9 | 45.89 |
| Lidar-UDA | [19] | 86.19 | 0.0 | 13.87 | 9.30 | 3.1 | 16.49 | 65.69 | 6.07 | 54.05 | 85.65 | 37.48 |
| CoSMix | [20] | 77.1 | 10.4 | 20.0 | 15.2 | 6.6 | 51.0 | 52.1 | 31.8 | 34.5 | 84.8 | 38.35 |
| SALUDA | [21] | 89.8 | 13.2 | 26.2 | 15.3 | 7.0 | 37.6 | 79.0 | 50.4 | 55.0 | 88.3 | 46.18 |
| Seg_3D_by_PC2D | (ours) | 92.06 | 0.02 | 27.13 | 15.45 | 8.28 | 43.35 | 78.72 | 52.59 | 71.31 | 90.42 | 47.93 |
| Supervised | (SK10SK10) | 91.32 | 35.78 | 11.83 | 0.03 | 20.17 | 22.87 | 90.55 | 79.60 | 63.93 | 88.01 | 50.41 |
| SK10NS (% IoU) | Car | Bicycle | Motorcycle | Truck | Other vehicle | Pedestrian | Driveable surf. | Sidewalk | Terrain | Vegetation | % mIoU | |
| Source-only | 1.79 | 0 | 0 | 0.01 | 0 | 0.5 | 6.65 | 0.07 | 0.05 | 0.26 | 0.93 | |
| T-UDA | [18] | 74.2 | 0.5 | 40.3 | 21.8 | 0.2 | 0.4 | 87.8 | 45.8 | 46.1 | 70.3 | 43.02 |
| Lidar-UDA | [19] | 73.48 | 0.85 | 15.86 | 0.9 | 25.67 | 40.78 | 87.44 | 42.31 | 47.88 | 83.22 | 43.64 |
| Seg_3D_by_PC2D | (ours) | 72.2 | 0 | 44.67 | 49.47 | 0 | 27.79 | 90.03 | 56.06 | 59.57 | 87.88 | 48.77 |
| Supervised | (NSNS) | 71.96 | 20.16 | 29.87 | 36.52 | 0 | 55.20 | 93.28 | 72.11 | 65.73 | 86.14 | 58.38 |
In this section we demonstrate the performance and capabilities of our method on Lidar Semantic Segmentation tasks. In particular, the 2D reliance of our method shows the most benefits when applied to domain adaptation. To that end, the multi-view projection pipeline is used to generate pseudo-labels for the target dataset, which can be used to fine-tune a pre-trained 3D model. This indirect approach ensures a fair comparison by using similar model architectures as the state-of-the-art, and allows for faster downstream inference. In addition, the results are improved compared to the direct output of the pipeline, for reasons discussed in section IV-D. First, we show the performance of our method in the Real-to-Real setting, where we compare it to the state-of-the-art methods in unsupervised domain adaptation. Then, we apply our method to the Simulation-to-Real setting, where we show that our approach can reach state-of-the-art performance on large, structured classes, although not reaching the state-of-the-art when averaged over all classes. Then, we demonstrate how our method can be used to segment rare classes, for which 3D annotations are not available, by only using 2D annotations for those classes and 3D annotations for other classes. Finally, we conduct an ablation study to demonstrate the importance of each component of our method.
IV-A Experimental setup
IV-A1 Datasets
Real-to-Real

### v3

@@ Abstract
3D semantic segmentation plays a pivotal role in autonomous driving and road infrastructure analysis, yet state-of-the-art 3D models are prone to severe domain shift when deployed across different datasets. In this paper, we propose an Unsupervised Domain Adaptation approach where a 3D segmentation model is trained on the target dataset using pseudo-labels generated by a novel multi-view projection framework. Our approach first aligns Lidar scans into coherent 3D scenes and renders them from multiple virtual camera poses to create large-scale synthetic 2D semantic segmentation datasets in various modalities. The generated datasets are used to train an ensemble of 2D segmentation models in point cloud view domain on each modality. During inference, the models process a large amount of views per scene; the resulting logits are back-projected to 3D with a depth-aware voting scheme to generate final point-wise labels. These labels are then used to fine-tune a 3D segmentation model in the target domain. We evaluate our approach Real-to-Real on the nuScenes and SemanticKITTI datasets. We also evaluate it Simulation-to-Real with the SynLidar dataset. Our contributions are a novel method that achieves state-of-the-art results in Real-to-Real Unsupervised Domain Adaptation, and we also demonstrate an application of our method to segment ‘rare classes’, for which target 3D annotations are not available, by only using 2D annotations for those classes and leveraging 3D annotations for other classes in a source domain.
1 Introduction
@@
3D semantic segmentation of Lidar scenes is crucial for applications such as autonomous driving (AD) and road infrastructure analysis. In particular, many applications require models to be able to generalize to unseen data, which may not be from the same domain as the training data, e.g. if the data was acquired in a different location or with a different sensor. The main approach to mitigate this problem is Unsupervised Domain Adaptation (UDA).
@@
Most existing methods are tailored for AD, thus making concessions on accuracy to guarantee real-time performance on single or few Lidar scans [24, 23, 18, 30].
@@
We propose a method that leverages aligned scenes to generate high quality pseudo-labels for the target domain, which are used to fine-tune a final 3D segmentation model pre-trained on the source domain. Our approach reaches state-of-the-art performance in UDA while introducing no computational overhead to the final 3D segmentation model. Our pseudo-label generation process belongs to the multi-view projection family of methods, where 3D scenes are analyzed via multiple 2D representations. We position virtual cameras around aligned Lidar scenes to generate rendered views (virtual camera images) from various perspectives. While similar methods use real camera images for training and inference [8], we create large-scale synthetic 2D semantic segmentation datasets by rendering both the 3D scenes and their corresponding ground truth annotations. This eliminates the need for additional camera sensors and allows to freely choose the optimal parameters for selecting and rendering the views.
We demonstrate the effectiveness of our method for UDA in Real-to-Real tasks on the nuScenes (NS) [2] and SemanticKITTI (SK) [1] datasets (NSSK and SKNS), and in a Simulation-to-Real task on the SynLidar (SL) [29] dataset (SLSK). We also demonstrate a potential application of our method to segment ‘rare classes’, where 3D annotations are not available for the desired target classes, but 2D annotations are available along with 3D annotations for other classes. Finally, we conduct a full ablation study to verify the importance of each component of our method. Our main contributions are the following:
@@
A method for Unsupervised Domain Adaptation in 3D semantic segmentation reaching state-of-the-art performance in
nuScenes SemanticKITTI and SemanticKITTI nuScenes
A novel multi-view projection framework for pseudo-label generation, with dataset generation capabilities
@@
A demonstration of an application of our method to segment ‘rare classes’ for which 3D annotations are not available
2 Related works
2.1 3D semantic segmentation
@@ 2 Related works
3D semantic segmentation is the task of assigning a semantic label to each point in a 3D point cloud. It is generally done either on Lidar point clouds [2, 1], or RGB-D images [6]. Methods are usually divided in 3D-based and 2D-based. 3D models directly operate on the point cloud using 3D sparse convolutions [5], while 2D models re-use 2D image segmentation methods by projecting the point cloud to a 2D plane [14]. Importantly, LiDAR-based 3D semantic segmentation models are particularly sensitive to domain shift due to sensor-specific sampling and sparsity effects, often more so than 2D vision models [24, 13].
@@
2D approaches generally project the point cloud to a single 2D image with channels representing various features, e.g. height, intensity, density, etc. Various projection approaches exist, including Bird’s Eye View (BEV) [14] and Range View (RV) [28]. Multi-view projection methods try to leverage 2D models even further by repeatedly using the same model on multiple views of the 3D scene.
2.2 Multi-view projection for segmentation
@@
P1.
Which projections to use for the views and with which inference model?
P2.
@@
Our proposed method uses a fast voting scheme on accumulated logits from a single 2D model.
In this paper, we focus on problem P1. Initially, [25] proposed multi-view projection for indoor RGB-D scenes, where the views are simply the RGB-D images themselves. For outdoor Lidar scenes, the same reasoning led [8] to use RGB camera images as views, with cameras calibrated to the Lidar [21]. However, this approach is limited as it requires using additional sensors, and the performance is tied to the quality of the sensor calibration. Mitigating this, [3] proposed to infer on rendered views of 3D scenes with an out-of-domain 2D model trained on camera images. This has the advantage of not requiring 3D labels or additional sensors, but comes with several limitations:
The domain gap between camera images and views of Lidar point clouds is significant, which negatively impacts 2D model performance.
The multi-view capability of the method is limited, as virtual camera placement must remain close to physically plausible viewpoints, e.g., those of real road users.
@@
We propose to solve these three limitations by leveraging 3D annotations. This makes our method more dependent on 3D annotated data, but allows it to achieve significantly better performance and handle dynamic classes. Closest to our approach, [12] proposes to use 3D annotations to generate 2D synthetic datasets to train a 2D model for downstream multi-view projection for indoor RGB-D scenes. Our method distinguishes itself in its application to outdoor Lidar scenes, where much additional processing is required to obtain satisfying views, e.g. scan aligning, as well as in the choices for the feature rendering and view selection. We propose an ensembling mechanism for handling the modality choice. We also demonstrate the potential of such methods for pseudo-label generation for unsupervised domain adaptation.
2.3 Unsupervised Domain Adaptation
UDA aims to adapt a model trained on a labeled source domain to perform well on an unlabeled target domain , where no ground-truth labels are available. Unlike supervised adaptation, only the source labels and unlabeled target samples can be exploited, while the evaluation is performed on a disjoint set . This setting is particularly relevant for autonomous driving, where annotated data is costly to acquire, but raw scans from the deployment domain can be collected at scale.
@@ 2.3 Unsupervised Domain Adaptation
Our method differs by relying entirely on a multi-view 2D projection pipeline to generate pseudo-labels for the target domain. These pseudo-labels are used to fine-tune a 3D segmentation model pretrained on the source domain. This approach reaches state-of-the-art performance in UDA.
3 Method
An overview of our method, which we call Seg3DbyPC2D , is shown in Fig. 1. To achieve UDA in 3D semantic segmentation, we simply pretrain a 3D model on the source dataset, then fine-tune it on the target dataset using pseudo-labels. Our main contribution is therefore our pseudo-label generation pipeline, which we describe in detail in this section.
In the following, we use the terms ‘scene’ to refer to a point cloud resulting from the accumulation of Lidar scans into a common frame, and ‘view’ to refer to a 2D image rendered from a 3D scene using a virtual camera.
3.1 Pseudo-label Generation
Our method can be decomposed in the source training phase and the target inference phase. In both phases, we obtain 3D scenes and sample virtual camera poses to render views of the scenes. In the source training phase, we use these rendered views, along with rendered views of the 3D segmentation masks, to generate synthetic segmentation datasets (PC2D) that will be used to train an ensemble of 2D models. In the target inference phase, the views are used for inference with the previously trained ensemble of models, yielding pseudo-labels for the target domain. The choices of parameters for Virtual Camera Placement (VCP) and view rendering from Multiple Modalities (MMods) are crucial and are detailed in the following subsections. To obtain final pseudo-labels in the target domain, our multi-view projection approach is based on [3], which we improve with occlusion handling.
3.1.1 Training with 3D annotations in 2D (PC2D)
A major contribution to our work compared to existing multi-view projection methods is the training of 2D models in rendered view domain. To that end, we generate large-scale synthetic datasets of 2D views of 3D scenes along with 2D segmentation masks, as shown in Fig. 2. We call these PC2D datasets.
@@ 3.1.1 Training with 3D annotations in 2D (PC2D)
Virtual camera poses are sampled arbitrarily around the 3D scene, which are then used to render 2D views. This way, we can generate an arbitrary number of (image, label) pairs, yielding large-scale datasets. Training on such datasets can be seen as a powerful data augmentation technique. The main limitation to this is that a limited number of annotated scenes will yield samples with less diversity as the total number of training samples increases.
@@
The use of PC2D datasets allows to train the models in a rendered view source domain, very close to the target inference domain, yielding greater freedom over the view generation steps compared to existing methods. For example, generating bird’s eye view depth maps is possible, while views in [3], are limited to ‘road user viewpoint’ rendering of Lidar intensity. The two following subsections describe how we choose the virtual camera placement and the rendering modalities to obtain the best performance.
3.1.2 Virtual Camera Placement
We propose to optimize the view generation by choosing relevant virtual camera poses.
@@ 3.1.2 Virtual Camera Placement
After experimentation, we retain 4 types of camera poses based on the original Lidar sensor poses, which are already describing the trajectory of the ego vehicle along the scene. An illustration of the 4 types of camera pose is shown in Fig. 2, in the VCP module, and we describe them below:
@@
carView: The camera is placed at a road user’s perspective, i.e. at the approximate height of a car (random yaw can be applied).
@@
conicView: The camera is placed on the basis of a large imaginary cone pointing downwards, with the apex at the Lidar sensor position. The camera points to the apex of the cone. This corresponds to a ‘drone’ or ‘lamp post’ view of the scene, and is also the closest to the views that human annotators would typically use to annotate a scene.
@@
topView: The camera is placed directly above the Lidar position, looking downwards, with variable height and roll. This corresponds to the usual ‘bird’s eye view’ (BEV) of the scene.
@@
bottomView: The camera is placed under the scene, looking upwards. This is the opposite of a BEV. Rendering Lidar point clouds from below allows to clearly see the colors of the flat surfaces of the ground without occlusions from the objects above.
3.1.3 Multi-Modalities
@@
Similar to camera pose generation, PC2D datasets allow us to choose the visualization modality for the training and inference images, i.e. the rendered colors. While multi-view projections methods based on RGB cameras are limited to RGB colors, we can choose to render modalities such as Lidar intensity, depth maps, and normals. The final projection and voting step does not limit the number of votes, we therefore propose to train multiple models on multiple modalities, and use them together in an ensemble. We use the following modalities:
For each modality, we generate a PC2D dataset, and train a separate model. We then use ensembles of models to process each view of the scene at inference time.
3.1.4 Projection and Fusion
@@
To obtain final 3D segmentation masks from 2D models’ outputs, we use a similar projection step as in [3], where 2D logits are back-projected to 3D and accumulated as votes. We propose adding occlusion to the projection step (OCL). Naive back-projection from 2D to 3D suffers from points being included in the projection even if they are occluded by objects. We use depth maps generated from meshes, themselves generated using [9] for occlusion. Using meshes has the advantage of allowing less sparse depth maps which is more appropriate for dense point clouds. Finally, we set a margin parameter to include points behind the depth map to account for the size of the occluding object. This final step yields the pseudo-labels for the target dataset.
4 Experiments
In this section we demonstrate the performance of our method for UDA on Lidar Semantic Segmentation tasks.
First, we show the performance of our method in both Real-to-Real and Simulation-to-Real settings. For both settings, we compare the results with the state-of-the-art methods in UDA. Then, we demonstrate how our method can be used to segment rare classes, for which 3D annotations are not available, by only using 2D annotations for those classes and 3D annotations for other classes. Finally, we conduct an ablation study to demonstrate the importance of each component of our method.
4.1 Datasets and Metrics